OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13)? 4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5

OpenStudy (shamil98):

m = y2-y1/x2-x1

OpenStudy (shamil98):

y - y1 = m ( x -x1)

OpenStudy (shamil98):

plug in the values !

OpenStudy (anonymous):

is it x − 4y = 5

OpenStudy (shamil98):

no, the slope is -4, so the answer would be C. x +4 y = 5 4y = -x + 5 y = -x/4 + 5